# FUSION: Forecast-Embedded Agent Scheduling with Service Incentive Optimization over Distributed Air-Ground Edge Networks

### v2

@@ B-E Economic Properties of Off-AIC2
The individual rationality of Off-AIC2 is guaranteed by the core design of its allocation (Step 2) and pricing (Step 3) procedures in Alg. 3, under the single-parameter assumption that each ES $s_{i}$ has a per-unit valuation and each AP $v_{k}$ has a per-unit cost.
@@
(i) Phase 1 (pivotal-rank selection). In Step 1 of Alg. 3, ESs are sorted in nonincreasing order of their average bids $\overline{\mathrm{bid}}_{i}$, and APs are sorted in nondecreasing order of their asks $\mathrm{ask}_{k}$. The pivotal ranks $(I^{\ast},K^{\ast})$ are chosen such that
@@
Alg. 3 only considers the top-$I^{\ast}$ ESs and top-$K^{\ast}$ APs for matching. Hence, for any winning pair $(i,k)$ with $x^{\mathrm{Off}}_{i,k}=1$, we have
@@
(ii) Phase 3 (critical pricing). Step 3 of Alg. 3 determines the actual payments within this feasible band using critical-value binary searches:
@@ Appendix C Properties of the Ordinal Potential
The proof follows a standard single-user deviation argument and relies on three ingredients: (i) separability of the non-congestion terms $S_{\mathbbm{j}}(\cdot)$, (ii) an integral–pointwise comparison that links the pointwise congestion penalty in user utilities with the integrated congestion cost in $\Phi(\bm{\pi})$, and (iii) monotonicity and convexity of the SP-side cost $K_{\mathbbm{i}}(\cdot)$. Full details are deferred to Appendices E-CE-G.
@@
Since each SD possesses a finite action set, any sequence of unilateral best-response (or better-response) updates generates a strictly increasing trajectory of the potential function $\Phi(\bm{\pi})$ and therefore converges in finitely many iterations to a pure-strategy NE. Equivalently, every local maximizer of $\Phi(\bm{\pi})$ corresponds to a pure-strategy NE of the online scheduling game. A formal proof of this finite-improvement property is provided in Appx. E-H.
Appendix D Detailed pseudo-code of PG-BRD
In this section, we provide the detailed pseudo-code of PG-BRD in Alg. 4.
2[2pt]
4[2pt]
5while not $\mathrm{converged}$ and $\mathbbm{t}<T_{\max}$ do
6
7 $\mathrm{converged}\leftarrow\textbf{true}$;
9
10 for each SD $u^{\mathrm{On}}_{\mathbbm{j}}$ in $\mathcal{P}$ do
11     let $a^{\mathrm{cur}}_{\mathbbm{j}}\leftarrow\pi_{\mathbbm{j}}$;
13
15
17
18    for each $a\in\mathcal{A}^{\mathrm{feas}}_{\mathbbm{j}}$ do
20
21
23
25        $\pi_{\mathbbm{j}}\leftarrow a^{\mathrm{best}}_{\mathbbm{j}}$;
26        $\mathrm{converged}\leftarrow\textbf{false}$;
27    else
28        $\pi_{\mathbbm{j}}\leftarrow a^{\mathrm{cur}}_{\mathbbm{j}}$;
29
30
31 $\mathbbm{t}\leftarrow\mathbbm{t}+1$;
32
Algorithm 4 Proposed PG-BRD
Appendix E Proof Details for the Ordinal Potential and PG-BRD
@@
E-A Matrix Formulation Based on the Global Assignment Matrix
@@
E-B Integral-Pointwise Sandwich Inequality
@@
E-C Single-User Deviation: Utility and Potential Changes
@@
Payment terms are absent from $\Delta\Phi$ by construction (see Appx. E-I).
E-D Approximate Linearization of Soft-Deadline Utility
@@
E-E Sign Consistency of Congestion Terms
@@
E-F Monotonicity and Convexity of SP Cost
@@
E-G Proof Sketch of Proposition 7
@@
  • By Appx. E-F, the SP-cost variation responds monotonically and convexly to the direction of load shift, further aligning the sign of $\Delta\Phi$ with that of $\Delta U_{\mathbbm{j}}$.
  • Payment differences $(p_{\mathbbm{j},\mathbbm{i}^{\prime}}-p_{\mathbbm{j},\mathbbm{i}})$ only redistribute welfare and are excluded from $\Delta\Phi$ (Appx. E-I), so they do not affect the ordinal comparison.
@@
E-H Finite-Improvement Property and Existence of NE
@@
E-I Treatment of Monetary Payments in the Potential Function

### v3

@@ B.2 Economic Properties of Off-AIC2
The individual rationality of Off-AIC2 is guaranteed by the core design of its allocation (Step 2) and pricing (Step 3) procedures in Alg. 2, under the single-parameter assumption that each ES $s_{i}$ has a per-unit valuation and each AP $v_{k}$ has a per-unit cost.
@@
(i) Phase 1 (pivotal-rank selection). In Step 1 of Alg. 2, ESs are sorted in nonincreasing order of their average bids $\overline{\mathrm{bid}}_{i}$, and APs are sorted in nondecreasing order of their asks $\mathrm{ask}_{k}$. The pivotal ranks $(I^{\ast},K^{\ast})$ are chosen such that
@@
Alg. 2 only considers the top-$I^{\ast}$ ESs and top-$K^{\ast}$ APs for matching. Hence, for any winning pair $(i,k)$ with $x^{\mathrm{Off}}_{i,k}=1$, we have
@@
(ii) Phase 3 (critical pricing). Step 3 of Alg. 2 determines the actual payments within this feasible band using critical-value binary searches:
@@ Appendix C Properties of the Ordinal Potential
The proof follows a standard single-user deviation argument and relies on three ingredients: (i) separability of the non-congestion terms $S_{\mathbbm{j}}(\cdot)$, (ii) an integral–pointwise comparison that links the pointwise congestion penalty in user utilities with the integrated congestion cost in $\Phi(\bm{\pi})$, and (iii) monotonicity and convexity of the SP-side cost $K_{\mathbbm{i}}(\cdot)$. Full details are deferred to Appendices D.3D.7.
@@
Since each SD possesses a finite action set, any sequence of unilateral best-response (or better-response) updates generates a strictly increasing trajectory of the potential function $\Phi(\bm{\pi})$ and therefore converges in finitely many iterations to a pure-strategy NE. Equivalently, every local maximizer of $\Phi(\bm{\pi})$ corresponds to a pure-strategy NE of the online scheduling game. A formal proof of this finite-improvement property is provided in Appx. D.8.
Appendix D Proof Details for the Ordinal Potential and PG-BRD
@@
D.1 Matrix Formulation Based on the Global Assignment Matrix
@@
D.2 Integral-Pointwise Sandwich Inequality
@@
D.3 Single-User Deviation: Utility and Potential Changes
@@
Payment terms are absent from $\Delta\Phi$ by construction (see Appx. D.9).
D.4 Approximate Linearization of Soft-Deadline Utility
@@
D.5 Sign Consistency of Congestion Terms
@@
D.6 Monotonicity and Convexity of SP Cost
@@
D.7 Proof Sketch of Proposition 7
@@
  • By Appx. D.6, the SP-cost variation responds monotonically and convexly to the direction of load shift, further aligning the sign of $\Delta\Phi$ with that of $\Delta U_{\mathbbm{j}}$.
  • Payment differences $(p_{\mathbbm{j},\mathbbm{i}^{\prime}}-p_{\mathbbm{j},\mathbbm{i}})$ only redistribute welfare and are excluded from $\Delta\Phi$ (Appx. D.9), so they do not affect the ordinal comparison.
@@
D.8 Finite-Improvement Property and Existence of NE
@@
D.9 Treatment of Monetary Payments in the Potential Function